OpenStudy (anonymous):

please help explain this? If the measure of θ is A radians and B degrees, then A=180B/π

OpenStudy (anonymous):

we can do this with knowing only two values for \(\theta\) and knowing converting from one measure to the other is linear we know \(0\) degrees is the same as \(0\) radians we know \(90\) degrees is \(\pi\) radians so to convert from degrees to radians, divide by 180 and multiply by \(\pi\) and to convert from radians to degrees, multiply by 180 and divide by \(\pi\) kind of like those "direct proportion" questions

OpenStudy (anonymous):

damn i mean " we know \(180\) degrees is \(\pi\) radians "

OpenStudy (anonymous):

sorry about that typo

OpenStudy (anonymous):

no problem :), so 0°=0 radians and 180°=π radians so what do i divide by 180 and multiply by π?

OpenStudy (anonymous):

sorry i just realized it was only the explanation i needed to do and not solve a problem :) thanks so much!!

OpenStudy (anonymous):

yw